OpenStudy (anonymous):

find the sum of the geometric series with a1=7, r=2, and n=14

OpenStudy (anonymous):

7(2^14 -1)/2^14 -1. The answer is 7.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u help me on these one find a1 for the geometric series with r=3 and S6=364

OpenStudy (anonymous):

your a is 364

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

just put the values given to you in the formula of sum of geometric progression

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

sn = a1 (1 - r^n) / (1 - r) a1 = 7 r = 2 n = 14 now we sub s14 = 7 (1 - 2^14) / (1 - 2) s14 = 7 (1 - 16384) / -1 s14 = 7 (- 16383) / -1 s14 = -114681/-1 s14 = 114681 sum of all 14 terms is 114681

OpenStudy (anonymous):

i did that but it was wrong

OpenStudy (texaschic101):

did you get the same answer ? Is the problem written correctly ? Did they want the sum or just the 14th term ?

OpenStudy (anonymous):

oh but i think i put 114681/2 i made a mistake i subtracted wrong

OpenStudy (anonymous):

they wanted the sum

OpenStudy (texaschic101):

it is just 114681

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help me on these on Find r for the geometric series S5=484, a1=4, and a5=324